Please find the dimension of b. cos[b W/V] where W has the dimension of work and V has the dimension of COS velocity M*LIT O M.LA2/TA2

Please explain all aspects of the question, no idea what it's asking
Answer #1

the question is asking the dimension of b.

In Cos[b*W/V] , b*W/V will be dimensionless because it's equal to angle. and angle has no dimension

b*W/V should be dimensionless.

given

W=[ML2T−2]


V= [LT−1]

b*W/V = b*  [ML2T−2]/[LT−1] = dimentionless

b*W/V = b* [MLT−1 ] = dimentioless (! have been taken cause 1 have no dimention )

so b= 1/ [MLT−1 ]

b=T/[ML]

First option is correct.